### v5

@@ 3.1 KV Cache Structure
See Figure 3(b) for a depiction of the three components in yellow, green, and purple, respectively. Each component is initialized during prefill, with top-$K$ tensors appended to the sink and recent tensors after any padding. Let $L_{sr}=L_{\text{sink}}+L_{\text{recent}}$; then, the exact structure of the KV cache has two cases:
3.1.1 Case 1: $L<L_{sr}$
In this case, the key tensor cache is constructed during prefill as $k_{1},...,k_{L},k_{\text{pad}:L+1},..,k_{\text{pad}:L_{sr}},k_{\text{topk}:1},..,k_{\text{topk}:K}$. This structure is no different than that described in Section 2.3, as the Top-$K$ tokens $x_{\text{topk}:\cdot}$ are themselves padding tensors in this case. Accordingly, during decoding, the KV cache is updated in-place in the same manner as previously described in Section 2.3.
3.1.2 Case 2: $L\geq L_{sr}$
@@
3.2 Prefill
In order to reduce memory usage of the decoding stage, SnapStream applies SnapKV compression during prefill to compress an output KV cache along the sequence dimension. Adherent to the KV cache structure described in Section 3.1, we only apply SnapKV to compress the evicted tokens $k_{L_{\text{sink}}+1},...,k_{L-L_{\text{recent}}}$. The sink token KV cache, recent token KV cache, and compressed SnapKV cache are then concatenated and transferred to the decoding node.
In preparing a KV cache for StreamingLLM decoding, we make two modifications to StreamingLLM during prefill:
  • We apply StreamingLLM after any positional encodings are applied to query, key, and value projections. This means that we do not apply positional encodings to cached KVs during decoding and use the original relative positional encodings. Although this may affect the ability of the model to generate beyond its trained attention window size as claimed in Section 3.2 of the StreamingLLM manuscript xiao2024efficientstreaminglanguagemodels, we believe that this change is well-motivated, as most modern open models are now trained to handle context lengths up to 128k grattafiori2024llama3herdmodels; deepseekai2025deepseekv3technicalreport; yang2025qwen3technicalreport, which is more than sufficient for most use cases.
@@ 3.2 Prefill
3.2.1 StreamingLLM Ring Buffer
In this section, we motivate and describe the implementation of the StreamingLLM rolling KV cache window, which is implemented as a ring buffer. For a naïve implementation of StreamingLLM using a static KV cache size $L_{\text{stream}}$, generating beyond $L_{\text{stream}}$ tokens one at a time involves inefficient slicing of the sink tokens, recent token window, and top-$k$ tokens from the KV cache and concatenating them all together with the new keys and values (see Lines 24-32 of Listing 2 in the Appendix). When the rolling window is implemented as a ring buffer, we only need to scatter new keys and values into the appropriate index during decoding (see Figure 3(d)). This requires us to prepare the ring buffer structure during the prefill phase, which involves two gather calls for the sequence indices before and after the prefill length index. See Listing 3 in the Appendix for a PyTorch-style implementation of the ring buffer construction, and Figure 4 for a visualization of where the two gather calls are placed along the sequence dimension.
@@ 3.2.1 StreamingLLM Ring Buffer
3.2.2 Recomputing Attention for SnapKV
In implementing SnapKV, we generally follow the original procedure outlined in Section 2.4, except that eviction candidates in SnapStream exclude sink tokens. We define $Q_{\text{obs}}=[q_{L-L_{\text{obs}}},...,q_{L}]$ as the query projections of the most recent $L_{\text{obs}}$ tokens and $K_{\text{evict}}=[k_{L_{\text{sink}}+1},...,k_{L-L_{\text{recent}}}]$ as the key projections of the set of eviction candidates. Then, the top-$K$ indices $I$ are computed exactly as in Equations 1 and 2, except that $W\in\mathbb{R}^{B\times H\times L_{\text{obs}}\times L_{\text{evict}}}$ is defined as
@@ 3.2.2 Recomputing Attention for SnapKV
We note that FlashAttention dao2022flashattentionfastmemoryefficientexact and other optimized attention implementations generally do not give their users access to the $O(L^{2})$ attention matrix $QK^{\intercal}$, as much of their value proposition comes from avoiding the expensive procedure of writing the attention matrix from SRAM to HBM. This has led to difficulties implementing sparse attention and efficient KV cache methods in libraries like vLLM vllmstreamingllmsupport2023. To get around this issue, our implementation recomputes Equation 3 outside of the fused attention kernel during prefill (see Section 3.2); we find that this amount of recompute constitutes less than 3% of the total prefill latency, but future works may optimize this step to avoid this additional compute.
@@
3.3 Decoding
By constructing the KV cache as a ring buffer, the decoding stage of a SnapStream deployment remains almost exactly the same as the standard decoding stage using static graphs described in Section 2.3, except that the in-place KV updates are performed at the modified position $L_{rb}$ defined in Equation 4. This means that we do not have to introduce any additional buffers from tensor indexing and concatenation that may be introduced by a naïve implementation of StreamingLLM (see Listing 2 in the Appendix). In particular, any fused kernels developed for decoding, such as those described in Section 4.1, may be used without modification.
$L_{rb}=(L+1-L_{sr})\mod L_{\text{recent}}+(L_{s}r)$ (4)
4 Implementation
Our goal in formulating SnapStream is to bridge the implementation gap between KV cache compression and LLM deployments, particularly for frameworks that use static graphs. We demonstrate the effectiveness of our approach by applying SnapStream to a real deployment of a state-of-the-art (SoTA) reasoning LLM, DeepSeek-R1-0528 deepseekr1, on the SN40L accelerator. In this section, we describe the kernel fusion and tensor sharding decisions, known as a “mapping”, for a single decoder layer of DeepSeek-R1-0528 on a single SN40L-16 node. Separate mappings are created for prefill and decoding that optimize for each phase’s respective SLOs. We also show how SnapStream integrates seamlessly into these mappings.
4.1 Prefill
@@
The minimum prefill deployment unit consists of a single SN40L-16 node executing with 16-way tensor parallelism (TP16). Prefill runs at batch size 1 with continuous batching. DeepSeek-R1-0528 consists of two types of decoder layers: dense FFN and sparse Mixture-of-Experts (MoE) layers. Figure 5 shows a high-level block diagram of the MoE prefill graph, including operators for SnapStream compression. The prefill graph is split into multiple fused kernels, with green boxes indicating kernel boundaries. The use of dataflow, large on-chip SRAM buffers, and a programmable interconnect enables unprecedented levels of operator fusion and pipelining. This results in increased operational intensity and reduced HBM traffic.
@@
  • StreamingLLM Ring Buffer: this kernel fuses four gather operations for Range 1 and Range 2 from Figure 4 for each of the key and value caches, as well as indexing for the sink tensors. All operations are equally sharded across the number attention heads $H$ into 16 sockets.
@@
      the recomputed $QK^{\intercal}$ GEMM described in Section 3.2.2
@@
      two gather operations using the top-$K$ indices $I$ for each of the key and value tensors
@@
Following the QKV, StreamingLLM Ring Buffer and SnapKV compression kernels, the MHA kernel implements multi-head attention using the QKV projections of the previous sections. Similarly to FlashAttention, we fuse the entire $QK^{\intercal}$ multiplication, softmax and $PV$ multiplication into a single kernel to avoid materializing the full attention matrix $P$. We distribute all operators in the MHA kernel along the number of attention heads $H$ equally across 16 sockets, and tile $K$ and $V$ into blocks of 8192 within each socket.
@@
Figure 6 illustrates a spatially mapped implementation of MoE FFN on SN40. Using spatial dataflow, we execute operators as a coarse-grained pipeline, where tensors are divided into tiles and streamed through sequential operators. Using TP16, each expert in a socket holds one-sixteenth of the weight matrix, requiring peer-to-peer (P2P) communication to gather activation fragments from all sockets before performing the down GEMM. The fused, pipelined, and tiled implementation enables P2P transfers to begin as soon as the first tile computation completes, effectively overlapping communication and computation, and further improving efficiency.
@@
4.2 Decoding
The decoding stage runs on a single SN40L-16 node. KV cache reuse significantly reduces the amount of computation required for decoding, effectively reducing the sequence dimension of most operations to 1. This allows us to compile even larger fused kernels for this stage, such that the entire computation graph is contained within two kernels: (a) a Multi-Head Latent Attention (MLA) kernel and (b) a Feed-Forward Network (FFN) kernel, as illustrated in Figure 7.
The MLA kernel includes QKV projections, attention computation, output projection, and the MoE router GEMM. The FFN section contains the shared and routed expert GEMMs. Both sections are implemented using fusion and spatial dataflow, as shown in Figure 6.
@@ 4.2 Decoding
The FFN section is executed under TP16, allowing weight reuse across all batch samples. Each socket computes all samples, but only on 1/16th of the routed expert slice, and only stores 1/16th of the expert weights locally. This enables grouping of tokens by expert, where all tokens requiring the same expert are processed together. During decoding, we employ the batch–sample fusion technique  nandkar2025speculativedecoding, where samples assigned to the same expert are fused and assigned to adjacent PCU systolic stages to increase FLOP utilization. This approach ensures that the weights for each expert are loaded only once across all 16 sockets and that the hardware achieves high compute efficiency through reuse and fused execution.
@@ 5.1 Accuracy
Benchmark
Llama 3.1
8B Instr.
+ StreamingLLM
8k
+ SAGE-KV
8k
+ SnapStream
8k
@@
Table 1: Benchmark results for Llama 3.1 8B Instruct on long sequence tasks.
We further evaluate the efficacy of SnapStream in a real deployment scenario with the mapping of DeepSeek-R1 described in Section 4 on a SN40L-16 node. We extend the SnapStream maximum KV cache length to 32k and evaluate the deployment on an improved long sequence benchmark, LongBench-v2, and two short-input reasoning tasks, AIME24 and LiveCodeBench. The results in Table 2 show that this real-life SnapStream deployment is able to maintain accuracy with respect to the full KV cache baselines, and in particular, SnapStream is able to outperform a similar deployment of fixed 32k sequence length without SnapStream on LongBench-v2.
@@
Table 2: Benchmark results for DeepSeek-R1-0528 on long sequence and reasoning tasks. “Full” and “32k” refer to full attention baselines with maximum sequence lengths of 128k and 32k respectively. “SnapStream-32k” is run with a maximum sequence length of 128k and a fixed 32k KV cache length.
@@
Table 3: Mean input/output token lengths for accuracy benchmarks on DeepSeek-R1-0528. Completions are sampled with greedy decoding and a maximum sequence length of 128k.
5.2 Performance
Table 4: Normalized latency breakdown per kernel for one prefill MoE layer of DeepSeek with input sequence size of 128K.
@@ 5.2 Performance
Table 5: Effect of SnapStream on maximum decode batch size and throughput for different prefill sequence sizes of a deployment of DeepSeek-R1-0528 on $16\timesSN$40L.
Prefill
Seq. Size
Compressed
Seq. Size
Max Batch Size
(No SnapStream)
Max Batch Size
Throughput
(No SnapStream)
Throughput
Throughput
Improvement
@@
In this section, we evaluate the impact of SnapStream on memory usage, maximum batch size, and overall decoding throughput. All experiments are conducted with DeepSeek-R1-0528 under the three KV cache compression configurations shown in Table 5. As described in Section 3.2, SnapKV compression is applied during the prefill stage. The additional compression logic introduces a modest latency overhead of approximately 2–5% of the total prefill time in the evaluated configurations. Table 4 presents the normalized latency breakdown for one MoE layer during prefill for DeepSeek-R1 with input sequence size of 128k. As expected, we find that the MHA and MoE FFN kernels dominate TTFT, and that the two additional SnapStream kernels contribute to $3\%$ of total latency.
@@
SnapStream compression enables larger decoding batch sizes within the fixed memory budget of a 16-socket node, assuming the model runs entirely in HBM. For long sequence sizes, the KV-cache contributes significantly to the memory footprint; to avoid replicating the KV-cache across sockets, we run the MLA operator with DP16. Combined with the $4\times$ compressed KV-cache, this results in a $4\times$ increase in the maximum attainable batch size, as summarized in Table 5.
@@
Table 5 also shows the impact of KV-cache compression on decoding throughput when using the maximum batch size allowed for each sequence length. Throughput was measured on a single SN40L-16 node and averaged over 10 runs. The reported values are scaled by 2.4 to account for the average number of tokens accepted per forward pass, assuming an 80% acceptance rate in Multi-Token Prediction (MTP). The combination of a smaller memory footprint and higher batch-level parallelism yields substantial end-to-end throughput gains — up to $4.3\times$ — for the evaluated configurations. These results highlight that SnapStream not only enables longer sequences and larger batches within the same memory budget but also directly improves decoding performance through better hardware utilization and reduced memory bandwidth pressure.
6 Related Work

### v6

@@ III-A KV Cache Structure
See Fig. 3(b) for a depiction of the three components in yellow, green, and purple, respectively. Each component is initialized during prefill, with top-$K$ tensors appended to the sink and recent tensors after any padding. Let $L_{sr}=L_{\text{sink}}+L_{\text{recent}}$; then, the exact structure of the KV cache has two cases:
III-A1 Case 1: $L<L_{sr}$
In this case, the key tensor cache is constructed during prefill as $k_{1},...,k_{L},k_{\text{pad}:L+1},..,k_{\text{pad}:L_{sr}},k_{\text{topk}:1},..,k_{\text{topk}:K}$. This structure is no different than that described in Section II-C, as the Top-$K$ tokens $x_{\text{topk}:\cdot}$ are themselves padding tensors in this case. Accordingly, during decoding, the KV cache is updated in-place in the same manner as previously described in Section II-C.
III-A2 Case 2: $L\geq L_{sr}$
@@
III-B Prefill
In order to reduce memory usage of the decoding stage, SnapStream applies SnapKV compression during prefill to compress an output KV cache along the sequence dimension. Adherent to the KV cache structure described in Section III-A, we only apply SnapKV to compress the evicted tokens $k_{L_{\text{sink}}+1},...,k_{L-L_{\text{recent}}}$. The sink token KV cache, recent token KV cache, and compressed SnapKV cache are then concatenated and transferred to the decoding node.
In preparing a KV cache for StreamingLLM decoding, we make two modifications to vanilla StreamingLLM during prefill:
  • We apply StreamingLLM after any positional encodings are applied to query, key, and value projections. This means that we do not apply positional encodings to cached KVs during decoding and use the original relative positional encodings. Although this may affect the ability of the model to generate beyond its trained attention window size as claimed in Section 3.2 of the StreamingLLM manuscript [32], we believe that this change is well-motivated, as most modern open models are now trained to handle context lengths up to 128k [14, 11, 33], which is more than sufficient for most use cases.
@@ III-B Prefill
III-B1 StreamingLLM Ring Buffer
In this section, we motivate and describe the implementation of the StreamingLLM rolling KV cache window, which is implemented as a ring buffer. For a naïve implementation of StreamingLLM using a static KV cache size $L_{\text{stream}}$, generating beyond $L_{\text{stream}}$ tokens one at a time involves inefficient slicing of the sink tokens, recent token window, and top-$k$ tokens from the KV cache and concatenating them all together with the new keys and values (see Lines 24-32 of Listing 2 in the Appendix). When the rolling window is implemented as a ring buffer, we only need to scatter new keys and values into the appropriate index during decoding (see Fig. 3(d)). This requires us to prepare the ring buffer structure during the prefill phase, which involves two gather calls for the sequence indices before and after the prefill length index. See Listing 3 in the Appendix for a PyTorch-style implementation of the ring buffer construction, and Fig. 4 for a visualization of where the two gather calls are placed along the sequence dimension.
@@ III-B1 StreamingLLM Ring Buffer
III-B2 Recomputing Attention for SnapKV
In implementing SnapKV, we generally follow the original procedure outlined in Section II-D, except that eviction candidates in SnapStream exclude sink tokens. We define $Q_{\text{obs}}=[q_{L-L_{\text{obs}}},...,q_{L}]$ as the query projections of the most recent $L_{\text{obs}}$ tokens and $K_{\text{evict}}=[k_{L_{\text{sink}}+1},...,k_{L-L_{\text{recent}}}]$ as the key projections of the set of eviction candidates. Then, the top-$K$ indices II are computed exactly as in Equations 1 and 2, except that $W\in\mathbb{R}^{B\times H\times L_{\text{obs}}\times L_{\text{evict}}}$ is defined as
@@ III-B2 Recomputing Attention for SnapKV
We note that FlashAttention [8] and other optimized attention implementations generally do not give their users access to the $O(L^{2})$ attention matrix $QK^{\intercal}$, as much of their value proposition comes from avoiding the expensive procedure of writing the attention matrix from SRAM to HBM. This has led to difficulties implementing sparse attention and efficient KV cache methods in libraries like vLLM [6]. To get around this issue, our implementation recomputes Equation 3 outside of the fused attention kernel during prefill (see Section III-B); we find that this amount of recompute constitutes less than 3% of the total prefill latency, but future works may optimize this step to avoid this additional compute.
@@
III-C Decoding
By constructing the KV cache as a ring buffer, the decoding stage of a SnapStream deployment remains almost exactly the same as the standard decoding stage using static graphs described in Section II-C, except that the in-place KV updates are performed at the modified position $L_{rb}$ defined in Equation 4. This means that we do not have to introduce any additional buffers from tensor indexing and concatenation that may be introduced by a naïve implementation of StreamingLLM (see Listing 2 in the Appendix). In particular, any fused kernels developed for decoding, such as those described in Section IV-A, may be used without modification.
$L_{rb}=(L-L_{\text{sink}}+L_{\text{recent}})\mod L_{\text{recent}}+L_{\text{sink}}$ (4)
IV Implementation
Our goal in formulating SnapStream is to bridge the implementation gap between KV cache compression and LLM deployments, particularly for frameworks that use static graphs. We demonstrate the effectiveness of our approach by applying SnapStream to a real deployment of a state-of-the-art (SoTA) reasoning LLM, DeepSeek-R1-0528 [9], on the SN40L accelerator. In this section, we describe the kernel fusion and tensor sharding decisions, known as a “mapping”, for a single decoder layer of DeepSeek-R1-0528 on a single SN40L-16 node. Separate mappings are created for prefill and decoding that optimize for each phase’s respective SLOs. We also show how SnapStream integrates seamlessly into these mappings.
IV-A Prefill
@@
The minimum prefill deployment unit consists of a single SN40L-16 node executing with 16-way tensor parallelism (TP16). Prefill runs at batch size 1 with continuous batching. DeepSeek-R1-0528 consists of two types of decoder layers: dense FFN and sparse Mixture-of-Experts (MoE) layers. Fig. 5 shows a high-level block diagram of the MoE prefill graph, including operators for SnapStream compression. The prefill graph is split into multiple fused kernels, with green boxes indicating kernel boundaries. The use of dataflow, large on-chip SRAM buffers, and a programmable interconnect enables unprecedented levels of operator fusion and pipelining. This results in increased operational intensity and reduced HBM traffic.
@@
  • StreamingLLM Ring Buffer: this kernel fuses four gather operations for Range 1 and Range 2 from Fig. 4 for each of the key and value caches, as well as indexing for the sink tensors. All operations are equally sharded across the number attention heads $H$ into 16 sockets.
@@
      the recomputed $QK^{\intercal}$ GEMM described in Section III-B2
@@
      two gather operations using the top-$K$ indices II for each of the key and value tensors
@@
Following the QKV, StreamingLLM Ring Buffer and SnapKV Compression kernels, the MHA kernel implements multi-head attention using the QKV projections of the previous sections. Similarly to FlashAttention, we fuse the entire $QK^{\intercal}$ multiplication, softmax and PVPV multiplication into a single kernel to avoid materializing the full attention matrix $P$. We distribute all operators in the MHA kernel along the number of attention heads $H$ equally across 16 sockets, and tile $K$ and $V$ into blocks of 8192 within each socket.
@@
Fig. 6 illustrates a spatially mapped implementation of the MoE FFN on SN40. Using spatial dataflow, we execute operators as a coarse-grained pipeline, where tensors are divided into tiles and streamed through sequential operators. Using TP16, each expert in a socket holds one-sixteenth of the weight matrix, requiring peer-to-peer (P2P) communication to gather activation fragments from all sockets before performing the down GEMM. The fused, pipelined, and tiled implementation enables P2P transfers to begin as soon as the first tile computation completes, effectively overlapping communication and computation, and further improving efficiency.
@@
IV-B Decoding
The decoding stage runs on a single SN40L-16 node. KV cache reuse significantly reduces the amount of computation required for decoding, effectively reducing the sequence dimension of most operations to 1. This allows us to compile even larger fused kernels for this stage, such that the entire computation graph is contained within two kernels: (a) a Multi-Head Latent Attention (MLA) kernel and (b) a Feed-Forward Network (FFN) kernel, as illustrated in Fig. 7.
The MLA kernel includes QKV projections, attention computation, output projection, and the MoE router GEMM. The FFN section contains the shared and routed expert GEMMs. Both sections are implemented using fusion and spatial dataflow, as shown in Fig. 6.
@@ IV-B Decoding
The FFN section is executed under TP16, allowing weight reuse across all batch samples. Each socket computes all samples, but only on 1/16th of the routed expert slice, and only stores 1/16th of the expert weights locally. This enables grouping of tokens by expert, where all tokens requiring the same expert are processed together. During decoding, we employ the batch–sample fusion technique  [21], where samples assigned to the same expert are fused and assigned to adjacent PCU systolic stages to increase FLOP utilization. This approach ensures that the weights for each expert are loaded only once across all 16 sockets and that the hardware achieves high compute efficiency through reuse and fused execution.
@@ V-A Accuracy
TABLE I: Benchmark results for Llama 3.1 8B Instruct on long sequence tasks.
Benchmark Llama 3.1 8B Instr. + StreamingLLM 8k + SAGE-KV 8k + SnapStream 8k
@@
We further evaluate the efficacy of SnapStream in a real deployment scenario with the mapping of DeepSeek-R1 described in Section IV on a SN40L-16 node. We extend the SnapStream maximum KV cache length to 32k and evaluate the deployment on an improved long sequence benchmark, LongBench-v2, and two short-input reasoning tasks, AIME24 and LiveCodeBench. The results in Table II show that this real-life SnapStream deployment is able to maintain accuracy with respect to the full KV cache baselines, and in particular, SnapStream is able to outperform a similar deployment of fixed 32k sequence length without SnapStream on LongBench-v2.
TABLE II: Benchmark results for DeepSeek-R1-0528 on long sequence and reasoning tasks. “Full” and “32k” refer to full attention baselines with maximum sequence lengths of 128k and 32k respectively. “SnapStream-32k” is run with a maximum sequence length of 128k and a fixed 32k KV cache length.
@@
TABLE III: Mean input/output token lengths for accuracy benchmarks on DeepSeek-R1-0528. Completions are sampled with greedy decoding and a maximum sequence length of 128k.
@@
V-B Performance
TABLE IV: Normalized latency breakdown per kernel for one prefill MoE layer of DeepSeek with input sequence size of 128K.
@@ V-B Performance
TABLE V: Effect of SnapStream on maximum decode batch size and throughput for different prefill sequence sizes of a deployment of DeepSeek-R1-0528 on a SN40L-16 node.
Prefill Seq. Size Compressed Seq. Size Max Batch Size (No SnapStream) Max Batch Size (SnapStream) Throughput (No SnapStream) Throughput (SnapStream) Throughput Improvement
@@
In this section, we evaluate the impact of SnapStream on memory usage, maximum batch size, and overall decoding throughput. All experiments are conducted with DeepSeek-R1-0528 under the three KV cache compression configurations shown in Table V. As described in Section III-B, SnapKV compression is applied during the prefill stage. The additional compression logic introduces a modest latency overhead of approximately 2–5% of the total prefill time in the evaluated configurations. Table IV presents the normalized latency breakdown for one MoE layer during prefill for DeepSeek-R1 with input sequence size of 128k. As expected, we find that the MHA and MoE FFN kernels dominate TTFT, and that the two additional SnapStream kernels contribute to $3\%$ of total latency.
@@
SnapStream compression enables larger decoding batch sizes within the fixed memory budget of a 16-socket node, assuming the model runs entirely in HBM. For long sequence sizes, the KV-cache contributes significantly to the memory footprint; to avoid replicating the KV-cache across sockets, we run the MLA operator with DP16. Combined with the $4\times$ compressed KV-cache, this results in a $4\times$ increase in the maximum attainable batch size, as summarized in Table V.
@@
Table V also shows the impact of KV-cache compression on decoding throughput when using the maximum batch size allowed for each sequence length. Throughput was measured on a single SN40L-16 node and averaged over 10 runs. The reported values are scaled by 2.4 to account for the average number of tokens accepted per forward pass, assuming an 80% acceptance rate in Multi-Token Prediction (MTP). The combination of a smaller memory footprint and higher batch-level parallelism yields substantial end-to-end throughput gains — up to $4.3\times$ — for the evaluated configurations. These results highlight that SnapStream not only enables longer sequences and larger batches within the same memory budget but also directly improves decoding performance through better hardware utilization and reduced memory bandwidth pressure.
VI Related Work